how many moles of h2 gas should be mixed with each mol of I2 vapour in order to convert 90% of it into HI [K=50]
1
Expert's answer
2019-10-10T03:46:25-0400

"H_2+I_2<=>2HI"

Let "a" moles of "H_2" react with "1mol" of "I_2"

At equilibrium, after "90\%" dissociation of "H_2" "K=\frac{[HI]^2}{[H_2][I_2]}=\frac{(1.8a)^2}{0.1a\times(1-0.9a)}=50"

On solving the above equation we get "a=0.6459"

Hence moles of "H_2" required"=0.6459mol"